6x^2+11x-10=0
x^2=7x
(x-5)(x+2)=18

Answer by mananth(16949)

6x^2+11x-10
6x^2-15x+4x-10=0
3x(2x-5)+2(2x-5)=0
(2x-5)(3x+2)=0
x= 2/5 OR x=-2/3
...
x^2=7x
divide by x
x=7
...
(x-5)(x+2)=18
x^2-3x-10=18
x^2-3x-28=0
x^2-7x+4x-28=0
x(x-7)+4(x-7)=0
(x+4)(x-7)=0
x=-4, x=7
...
m.ananth@hotmail.ca